OpenStudy (anonymous):

Find the length of the straight line y = mx from x=0 to x=a, where m and a are positive constants.

OpenStudy (anonymous):

\[\sqrt{(a-0)^2 + (ma-0)^2}\] = distance from point 1 to point 2 = length of the line segment from x=0 to x=a

OpenStudy (anonymous):

Because the distance formula between two points is \[\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

Thank you! Is there a way to solve this via geometry as well?

OpenStudy (anonymous):

(Pythagorean theorem)

OpenStudy (anonymous):

Yes, by drawing a triangle, which is actually where this formula for distance comes from.

OpenStudy (anonymous):



OpenStudy (anonymous):

see?

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh wow thank you very much!!! This is so helpful

OpenStudy (anonymous):

My pleasure. :)